An iron cube of mass 480g and density 8g/cm³ is suspended by a string so that its half submerged into oil of density 0.8g/cm³ . Find tension on the string
1
Expert's answer
2019-10-16T09:31:21-0400
"F_{t} - mg+ pgV = 0"

Where Ft- is tension on the string, mg- gravity, pgV - archimedes force.


"F_{t} = mg - pg\frac{m}{2p_{iron}}"

Where p -is density of oil, p iron - density of iron cube


"F_{t} = 4.71 - 0.24 = 4.47 N"